A long, straight wire of length 0.82 m carries current I = 1.6 Ain a region where B = 2.4 T. If the force on the wire is 1.4 N, what is the smallest possible angle between the field and the wire?________ °
Answer #1

Solution)

We know,

F/l = I* B * sin(theta)

where theta = angle between field and wire

On substituting values we get,

1.4 N/0.82 m =1.6 A*2.4 T * sin theta

Theta= 26.39 degrees

===========